# A review on recent progresses in the study of low energy and scatterings

### v1

@@ Abstract
We summarize new results on the studies of and scatterings. They include the finding of a negative-parity nucleon pole with a mass lower than the nucleon mass, and the pole trajectory of with respect to varying pion masses. The results are obtained from model-independent dispersion analyses. We also study the thermal properties of based on the model and method.
@@
In recent few years considerable progress has been made in understanding low-energy QCD, through the study of and scatterings. In particular, a novel subthreshold pole has been found in the channel, with the pole position, MeV. This result is obtained by a Steiner or Roy–Steiner equation analysis [1], as inspired by an earlier analysis using the product representation for partial wave amplitudes [2]. This result is totally unexpected, since it sounds completely counterintuitive to imagine a resonant pole below the ‘lowest’ lying and hence stable nucleon pole.
@@
In addition to the sector, there are also new results on the / resonance. In a series of recent publications [3, 4, 5], we have achieved a better understanding of the property of this very important particle, since it is widely believed that plays a crucial role in the spontaneous chiral symmetry breaking mechanism. We will summarize our new results on in Section 3. Section 4 is dedicated to the conclusions and perspectives for future explorations, especially the thermal property of .
@@ 2 The (920) pole
The existence of pole was originally suggested in Ref. [2], where the authors analyzed all six low partial waves (two and four waves) for scattering amplitudes. According to the product representation [6, 7, 8, 9], different contributions to the scattering phase shifts are additive. Particularly, resonance pole and virtual state pole contributions are positive whereas bound state pole contribution and ‘left-hand cut’ contribution are negative. The left-hand cut or the background contribution is known in potential scattering theory to be [10], where is the channel momentum and the interaction range of the potential (exact for truncated potentials [11]). In relativistic scattering theory the background contribution is represented by an integral on non-unitary cuts. It is estimated using tree-level chiral amplitudes in Ref. [2], and amplitudes in Ref. [12], both of which are qualitatively in agreement with the potential scattering case. After all known contributions to the total phase shift are carefully taken into account, it is revealed that a large discrepancy with the experiment persists in the channel among all six partial waves 111 Actually, a significant mismatch also arises in the channel. This is compensated by an accompanying virtual pole contribution, in addition to the nucleon bound state pole. The existence of such a virtual pole is mandated by the ‘elementariness’ of the nucleon [13, 14]. This discrepancy can only be accounted for by the existence of second-sheet poles. As such, a resonance pole located at MeV is needed to accommodate the experimental data. In addition, it is examined that an estimation of the left-hand cut contribution at level only slightly alters the pole position [2, 12]. See Fig.1 for an illustration. Furthermore, the pole has been subsequently confirmed in the -matrix [15] and method [16] unitarization approximation of the scattering amplitude. Nevertheless, it is worth pointing out that a -matrix unitarization approximation suffers some fundamental-level problems. There are three fundamental properties of the matrix: unitarity, analyticity, and crossing symmetry. The -matrix unitarization often violates the latter two [17, 18, 19] (for a pedagogical review, see also Ref. [20]). Hence, the -matrix unitarization method and its variations are generally not suitable for the discussion of distant poles.
@@
Nevertheless, much work remains to be done in understanding the physical properties of such a pole hidden far away from the physical region. Before proceeding, it is worth pointing out that the nucleon Regge trajectory as inspired by the generalized McDowell symmetry [25] appears to necessitate such a subthreshold pole, as can be seen in Fig. 3. Another way to probe the nature of is to study its pole trajectory with respect to varying pion masses, which could be useful for future lattice studies. In [26], using the linear -model and the method as the unitarization scheme, at tree level, it is found that the pole moves towards the -cut on the real axis in the -plane as pion mass increases (see Fig.4 for an illustration). It finally disappears from the second sheet and moves across the -cut at a certain point to the other sheet defined by the logarithmic branch points. This may not necessarily be the final destiny of the pole, however. In principle, when the pole approaches the real axis, there is no compelling reason it has to hit the -cut. For example, it might be possible that the poles fall at some point between the -cut and the right-hand cut on the real axis and then split into two virtual states. Further one-loop calculation is necessary and still in progress.
@@ 3 Dynamical properties of resonance in the case of unphysical pion masses and high temperatures
The existence of resonance was once a topic of debate for decades, until it was finally settled using the dispersion techniques [27, 28, 29] 333For a review, one is referred to, for example, Ref. [20]. Nevertheless, to the best of the authors’ knowledge, there has never been conclusive evidence that the resonance takes the pivotal role of the prominent meson in spontaneous chiral symmetry breaking, though it may already be rather widely embraced. 444See for example, Refs. [30]
@@
Recently, considerable progress has been made in this direction. In lattice QCD studies, more and more phase-shift data have been accumulated with different input pion masses. An investigation on the pole trajectory with respect to varying pion masses becomes available. In [31, 32], pole positions as extracted from lattice data using a simple -matrix method, are found to be a resonance pole at around MeV for MeV, and a bound state pole at around MeV for MeV. There are also some theoretical work using unitarization method such as IAM and the unitarized PT to discuss the resonance pole trajectories in scattering [33, 34, 35]. The result shows that as the pion mass increases, the pole moves toward the real axis and become two virtual poles. Then one virtual pole moves towards the threshold and across the threshold becoming a bound state, which qualitatively agrees with the lattice result. However, as previously pointed out, widely used matrix unitarizations often have severe shortcomings, which do not match the lattice data that are obtained from ab initio calculations. In addition, since crossing symmetry is widely believed to be important in determining the resonance, incorporating crossing symmetry in the analysis of lattice data is necessary. A first trial in this direction is in [36] where the PKU representation was utilized in determining the pole position from the lattice data, in which the left-hand cut contribution was estimated using the PT scattering results. The result shows that a virtual state is accompanying the bound state when MeV. However, since becomes a bound state, it will also generate a left-hand cut through the cross-channel bound-state exchange processes, which does not present in the estimation [37]. These unsatisfactory situations are resolved in [3], where modified Roy equations are constructed to accommodate the new situation of bound states for large pion masses, and the phase shifts obtained by numerically solving these equations agree well with the lattice data. It is found that at MeV, the pole moves to MeV and at MeV, the pole becomes a bound state at MeV. Besides the bound state found when MeV, there is no virtual pole but a subthreshold resonance pole located at MeV. All the singularities on the second sheet are inside the Lehmann ellipse as shown in Fig. 5. To understand this subthreshold resonance pole, it is argued that another right-moving virtual state (VS-II) appears from the branch point of the left-hand cut generated by the after it becomes a bound state when pion mass is large. The subthreshold resonance pole is generated after VS-II meets the left-moving virtual state pole (VS-I) originated from poles after they hits the real axis below the threshold when pion mass becomes large enough. The suggested pole trajectory is shown in Fig. 6.
@@
The success of the linear model certainly deserves more respects and attentions. A further investigation is to study the thermal properties of the pole (namely the resonance) in the model. Similar to the zero-temperature case, the cross-channel contribution to the amplitude would also affect the behavior of the poles in the finite temperature scattering amplitude. The method can be extended to the finite temperature case [5]. It is rather surprising to have found that thermal trajectory of the pole is qualitatively very similar to the trajectory with varying as shown in Fig. 7. For example, when MeV turns into a bound state from a virtual pole as well. 666Similar result is also obtained in [40, 41] using the leading order result of the model, but since the crossed channels are not included, there are no subthreshold poles generated from the branch point of the left-hand cut in these analyses. Also the thermal vacuum property is carefully investigated and it is found that in linear model the vacuum still remains stable at the moment when turns into a bound state [5, 4]. It is noticed that the pole not only turns into a bound state, but will also eventually approach the pion-pole position. This actually is not an accident, but dictated by the asymptotically restored symmetry ( in reality) at the high-temperature limit. It is worth noticing that such a restoration is not yet fulfilled in usual thermal PT calculations.
The realization of chiral symmetry becomes much subtler when nucleons are involved. The restoration of chiral symmetry at high temperatures, in the chiral limit, would require parity doubling of nucleons and one may wonder the role of may play here. The subtlety is that, if the nucleon became massless in the chiral symmetric phase, there is no need to introduce a new parity partner for the nucleon, because a massless fermion can be decomposed into two helicity mass eigenstates with opposite parity (see for example [42]). On the other hand, a nucleon does not need to be massless in the chiral symmetric phase. This can be achieved at the effective lagrangian level by explicitly introducing the parity partner of the nucleon [43, 44, 45]. The chiral partner of the nucleon is usually expected to be (see for example Ref. [46, 47]). In lattice studies of thermal QCD, there is evidence that the becomes degenerate with the nucleon at high temperatures[48, 49]. What role the pole may play in this process would become an interesting question.

### v2

# Review on recent progress in the study of the subthreshold singularity and the meson

@@ Abstract
We summarize recent results on studies of and scatterings. They include the finding of a negative-parity nucleon pole with a mass lower than the nucleon mass, and the pole trajectory of as the pion mass varies. The results are obtained from model-independent dispersion analyses. We also study the thermal properties of based on the model and method.
@@
In recent years, considerable progress has been made in understanding low-energy QCD, through the study of and scatterings. In particular, a novel subthreshold pole has been found in the channel, with the pole position, MeV. This result is obtained by a Steiner or Roy–Steiner equation analysis [1], as inspired by an earlier analysis using the product representation for partial wave amplitudes [2]. This result is totally unexpected, since it sounds completely counterintuitive to imagine a resonant pole below the ‘lowest’ lying and hence stable nucleon pole.
@@
In addition to the sector, there are also new results on the / resonance. There has been a long history in the study of this particle. Due to page limitation here we only refer to earlier review papers, such as Refs. [3, 4, 5] for previous developments in this area. In a series of recent publications [6, 7, 8], we have achieved a better understanding of the property of this very important particle, since it is widely believed that plays a crucial role in the spontaneous chiral symmetry breaking mechanism. We will summarize our new results on in Section 3. Section 4 is dedicated to the conclusions and perspectives for future explorations, especially the thermal property of .
@@ 2 The (920) pole
The existence of pole was originally suggested in Ref. [2], where the authors analyzed all six low partial waves (two and four waves) for scattering amplitudes. According to the product representation [9, 10, 11, 12], different contributions to the scattering phase shifts are additive. Particularly, resonance pole and virtual state pole contributions are positive whereas bound state pole contribution and ‘left-hand cut’ contribution are negative. The left-hand cut or the background contribution is known in potential scattering theory to be [13], where is the channel momentum and the interaction range of the potential (exact for truncated potentials [14]). In relativistic scattering theory the background contribution is represented by an integral on non-unitary cuts. It is estimated using tree-level chiral amplitudes in Ref. [2], and amplitudes in Ref. [15], both of which are qualitatively in agreement with the potential scattering case. After all known contributions to the total phase shift are carefully taken into account, it is revealed that a large discrepancy with the experiment persists in the channel among all six partial waves 111Actually, a significant mismatch also arises in the channel. This is compensated by an accompanying virtual pole contribution, in addition to the nucleon bound state pole. The existence of such a virtual pole is mandated by the ‘elementariness’ of the nucleon [16, 17]. This discrepancy can only be accounted for by the existence of second-sheet poles. As such, a resonance pole located at MeV is needed to accommodate the experimental data. In addition, it is examined that an estimation of the left-hand cut contribution at level only slightly alters the pole position [2, 15]. See Fig.1 for an illustration. Furthermore, the pole has been subsequently confirmed in the -matrix [18] and method [19] unitarization approximation of the scattering amplitude.
Actually, there are even earlier analyses in the channel of scatterings, which suggest the possible existence of a pole with a mass around 1100 MeV (and a width around 400 MeV) which is very near to the threshold rather than below the nucleon mass [20, 21, 22]. Nevertheless, it is worth pointing out that a -matrix unitarization approximation suffers some fundamental-level problems. There are three fundamental properties of the matrix: unitarity, analyticity, and crossing symmetry. The -matrix unitarization often violates the latter two [23, 24, 25] (for a pedagogical review, see also Ref. [4]). Hence, the -matrix unitarization method and its variations are generally not suitable for the discussion of distant poles. Similar problems also exist in the other unitarization approaches such as Bethe-Salpeter equation, unitarized PT. As a result, though a similar near threshold pole was found in [22], the authors failed to recognize the pole as a genuine state but tended to interpret it as an effect that mimics the missing cut [26, 22]. A comment on Ref. [22] can be found in Ref. [19].
@@
Nevertheless, much work remains to be done in understanding the physical properties of such a pole hidden far away from the physical region. Before proceeding, it is worth pointing out that the nucleon Regge trajectory as inspired by the generalized McDowell symmetry [31] appears to necessitate such a subthreshold pole, as can be seen in Fig. 3. Another way to probe the nature of is to study its pole trajectory as the pion mass varies, which could be useful for future lattice studies. In [32], using the linear -model and the method as the unitarization scheme, at tree level, it is found that the pole moves towards the -cut on the real axis in the -plane as pion mass increases (see Fig.4 for an illustration). It finally disappears from the second sheet and moves across the -cut at a certain point to the other sheet defined by the logarithmic branch points. This may not necessarily be the final destiny of the pole, however. In principle, when the pole approaches the real axis, there is no compelling reason it has to hit the -cut. For example, it might be possible that the poles fall at some point between the -cut and the right-hand cut on the real axis and then split into two virtual states. Further one-loop calculation is necessary and still in progress.
@@ 3 Dynamical properties of resonance in the case of unphysical pion masses and high temperatures
The existence of resonance was once a topic of debate for decades, until it was finally settled using the dispersion techniques [33, 34, 35] 333For a review, one is referred to, for example, Ref. [4].. Nevertheless, to the best of the authors’ knowledge, there has never been conclusive evidence that the resonance takes the pivotal role of the prominent meson in spontaneous chiral symmetry breaking, though it may already be rather widely embraced 444See for example, Refs. [36].
@@
Recently, considerable progress has been made in this direction. In lattice QCD studies, more and more phase-shift data have been accumulated with different input pion masses. An investigation on the pole trajectory with respect to varying pion masses becomes available. In [37, 38], pole positions as extracted from lattice data using a simple -matrix method, are found to be a resonance pole at around MeV for MeV, and a bound state pole at around MeV for MeV. There are also some theoretical work using unitarization method such as IAM and the unitarized PT to discuss the resonance pole trajectories in scattering [39, 40, 41]. The result shows that as the pion mass increases, the pole moves toward the real axis and become two virtual poles. Then one virtual pole moves towards the threshold and across the threshold becoming a bound state, which qualitatively agrees with the lattice result. However, as previously pointed out, widely used matrix unitarizations often have severe shortcomings, which do not match the lattice data that are obtained from ab initio calculations. In addition, since crossing symmetry is widely believed to be important in determining the resonance, incorporating crossing symmetry in the analysis of lattice data is necessary. A first trial in this direction is in [42] where the PKU representation was utilized in determining the pole position from the lattice data, in which the left-hand cut contribution was estimated using the PT scattering results and the crossing symmetry was imposed by the BNR relations. The result shows that a virtual state is accompanying the bound state when MeV. However, since becomes a bound state, it will also generate a left-hand cut through the cross-channel bound-state exchange processes, which does not present in the estimation [43]. These unsatisfactory situations are resolved in [6], where modified Roy equations are constructed to accommodate the new situation of bound states for large pion masses, and the phase shifts obtained by numerically solving these equations agree well with the lattice data. It is found that at MeV, the pole moves to MeV and at MeV, the pole becomes a bound state at MeV. Besides the bound state found when MeV, there is no virtual pole but a subthreshold resonance pole located at MeV. All the singularities on the second sheet are inside the Lehmann ellipse as shown in Fig. 5. To understand this subthreshold resonance pole, it is argued that another right-moving virtual state (VS-II) appears from the branch point of the left-hand cut generated by the after it becomes a bound state when pion mass is large. The subthreshold resonance pole is generated after VS-II meets the left-moving virtual state pole (VS-I) originated from poles after they hits the real axis below the threshold when pion mass becomes large enough. The suggested pole trajectory is shown in Fig. 6.
@@
The success of the linear model certainly deserves more respects and attentions. A further investigation is to study the thermal properties of the pole (namely the resonance) in the model. Similar to the zero-temperature case, the cross-channel contribution to the amplitude would also affect the behavior of the poles in the finite temperature scattering amplitude. The method can be extended to the finite temperature case [8]. It is rather surprising to have found that thermal trajectory of the pole is qualitatively very similar to the trajectory with varying as shown in Fig. 7. For example, when MeV turns into a bound state from a virtual pole as well. 666Similar result is also obtained in [46, 47] using the leading order result of the model, but since the crossed channels are not included, there are no subthreshold poles generated from the branch point of the left-hand cut in these analyses. Also the thermal vacuum property is carefully investigated and it is found that in linear model the vacuum still remains stable at the moment when turns into a bound state [8, 7]. It is noticed that the pole not only turns into a bound state, but will also eventually approach the pion-pole position. This actually is not an accident, but dictated by the asymptotically restored symmetry ( in reality) at the high-temperature limit.
It is worth noticing that such a restoration is hard to be fulfilled in usual thermal PT calculations, since PT is constructed just in the broken phase. Nevertheless, there are also some hints of the chiral restoration from PT calculations. The chiral condensate is calculated to be decreasing with temperature [48], and the scalar and pseudoscalar susceptibility are estimated to be degenerate around a critical temperature within the unitarized PT [49, 50]. However, though the pole trajectory obtained using the unitarized PT shows a decreasing mass with increasing temperature and seems to display a trend towards the real axis at temperatures roughly above 150MeV [51, 52, 53] 777To produce this behavior the Adler zero must be incorporated into the IAM formalism, otherwise the behavior is quite different [54]., near the chiral restoration temperature, the pole would not turn to a bound state or tend to the pion mass, which demonstrates the limitation of unitary approximation based on PT.
The realization of chiral symmetry becomes much subtler when nucleons are involved. The restoration of chiral symmetry at high temperatures, in the chiral limit, would require parity doubling of nucleons and one may wonder the role of may play here. The subtlety is that, if the nucleon became massless in the chiral symmetric phase, there is no need to introduce a new parity partner for the nucleon, because a massless fermion can be decomposed into two helicity mass eigenstates with opposite parity (see for example [55]). On the other hand, a nucleon does not need to be massless in the chiral symmetric phase. This can be achieved at the effective lagrangian level by explicitly introducing the parity partner of the nucleon [56, 57, 58, 59]. The chiral partner of the nucleon is usually expected to be (see for example Ref. [60, 61]). In lattice studies of thermal QCD, there is evidence that the becomes degenerate with the nucleon at high temperatures [62, 63]. What role the pole may play in this process would become an interesting question.